Halliford Corporation expects to have earnings this coming year of

per share. Halliford plans to retain all of its earnings for the next two years. For the subsequent two years, the firm will retain of its earnings. It will then retain of its earnings from that point onward. Each year, retained earnings will be invested in new projects with an expected return of

per year. Any earnings that are not retained will be paid out as dividends. Assume Halliford's share count remains constant and all earnings growth comes from the investment of retained earnings. If Halliford's equity cost of capital is , what price would you estimate for Halliford stock?

Note: Remenber that growth rate is computed as: retention rate rate of return.
Answer #1

To be placed in chart below

18.89%* 45%= 0.085005 or 8.50%

18.89%%* 20%= 0.03778 or 3.78%

3.36*(1+18.89%)= 3.994704 or 3.99

3.99*(1+18.89%)= 4.743711 or $4.74

4.74*(1+8.50%)= 5.1429 or 5.14

5.14*(1+8.50%)= 5.5769 or $5.58

5.58*(1+3.78%)= 5.790924 or 5.79

1-100%= 0       1-45%= 0.55     1-20%= 0.8 or 80%     3.36*0%= $0.00

4.74* 0.55= 2.607 or 2.61

5.14* 55%= 2.827 or 2.83

5.58*80%= 4.464 or 4.46

5.29* 80%= 4.632 or 4.63

3.78%= 0.0378

Note: bottom is the formula, answer is 60.61 as shown

20210714_141250-min.jpg

answered by: Andrew San Andres